How do I find all the values of w satisfying the equation?
6-5/w+2=-4/w-1

Answer by checkley71(8403)

6-(5/W)+2=-(4/W)-1
-5/W+4/W=-6-2-1
(-5+4)/W=-9
-1/W=-9
-9W=-1
W=-1/-9
W=1/9